OpenStudy (anonymous):

Name the property of equality that justifies: If x + 3 = 5, then x = 2.

Directrix (directrix):

Subtraction Property of Equality From both sides of the equation x + 3 = 5, 3 was subtracted. That gave x +3 -3 = 5 - 3 which gives x = 2. You could also argue that Addition Property of Equality is the property if you observe that subtracting 3 is the same as adding negative 3.